# Long-time Freeness in the Kicked Top

### v1

@@ V Conclusions and Discussion
  • Large deviation theory of freeness. Guided by approaches in turbulence, chaos, and eigenstate delocalization, in this work, we introduced a large deviation theory of freeness, which shall apply generically to exponentially decaying $2n$-OTOCs. The associated freeness time-scale $\tau_{n}$ describe the time-dependent fluctuations of the spectrum of time-evolving observables and indicate a multifractal behavior if $\tau_{n}>\tau_{\rm typ}$ dependent on $n$, and a form of monofractality if $\tau_{n}=\tau_{\rm typ}$ for every $n$. In quantum systems with a few-body classical chaotic limits, the exponential decay of the $4$-OTOC was interpreted from the point of view of mixing and associated with Ruelle-Pollicot resonances by Refs.García-Mata et al. (2018); Notenson et al. (2023), see also Ref. Polchinski (2015). Within this framework, it would be intriguing to explore whether the mono/multifractal behaviors in the large deviation theory of freeness can be interpreted in terms of these resonances.
    Although our analysis stems from chaos in few-body semiclassical systems, it should extend to any context where freeness is achieved exponentially, including many-body systems. For example, recent findings report exponential decay in free cumulants of order $n$ within dual-unitary circuits Chen and Kudler-Flam (2024), suggesting promising directions for future exploration of the freeness time-scale in many-body settings.
@@
    For instance, we have shown that the leading-order ETH predictions fail to capture the early-time behavior of the square-commutator (see Sec. III.3) or the late-time fluctuations of the free cumulants (see Sec. III.1). This highlights the need to extend ETH to account for finite system size corrections.
    Finally, we explored the development of freeness and free-cumulant dynamics up to their $\mathcal{O}(1/D)$ saturation level. At even longer time scales, signatures of level repulsion may appear. This effect is known at the level of the two-point function Cotler et al. (2017b); Delacretaz (2020); Altland et al. (2021); Yoshimura et al. (2023); Bouverot-Dupuis et al. (2024), and it is also expected at higher orders Cotler et al. (2017a); Cotler and Hunter-Jones (2020). While this is consistent with asymptotic long-time freeness, it invites further investigation into how higher-order spectral fluctuations manifest in observable quantities.
@@ Acknowledgements
We thank X. Turkeshi, B. Pain and J. Kudler-Flam for their insightful discussions, L. Foini for valuable feedback on our work, and J. Kurchan for collaboration on related projects. We acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 -390534769, and DFG Collaborative Research Center (CRC) 183 Project No. 277101999 - project B02.
Appendix A Diagrammatic representation of
the Full Eigenstate Thermalization Hypothesis

### v2

@@ 5 Conclusions and Discussion
  • Large deviation theory of freeness. Guided by approaches in turbulence, chaos, and eigenstate delocalization, in this work, we introduced a large deviation theory of freeness, which shall apply generically to exponentially decaying $2n$-OTOCs. The associated freeness time-scale $\tau_{n}$ describe the time-dependent fluctuations of the spectrum of time-evolving observables and indicate a multifractal behavior if $\tau_{n}>\tau_{\rm typ}$ dependent on $n$, and a form of monofractality if $\tau_{n}=\tau_{\rm typ}$ for every $n$. In quantum systems with a few-body classical chaotic limits, the exponential decay of the $4$-OTOC was interpreted from the point of view of mixing and associated with Ruelle-Pollicot resonances by Refs.[109, 110], see also Ref. [112]. Within this framework, it would be intriguing to explore whether the mono/multifractal behaviors in the large deviation theory of freeness can be interpreted in terms of these resonances.
    Although our analysis stems from chaos in few-body semiclassical systems, it should extend to any context where freeness is achieved exponentially, including many-body systems. For example, recent findings report exponential decay in free cumulants of order $n$ within dual-unitary circuits [85], suggesting promising directions for future exploration of the freeness time-scale in many-body settings.
@@
    For instance, we have shown that the leading-order ETH predictions fail to capture the early-time behavior of the square-commutator (see Sec. 3.3) or the late-time fluctuations of the free cumulants (see Sec. 3.1). This highlights the need to extend ETH to account for finite system size corrections.
    Finally, we explored the development of freeness and free-cumulant dynamics up to their $\mathcal{O}(1/D)$ saturation level. At even longer time scales, signatures of level repulsion may appear. This effect is known at the level of the two-point function [113, 114, 115, 116, 117], and it is also expected at higher orders [33, 35]. While this is consistent with asymptotic long-time freeness, it invites further investigation into how higher-order spectral fluctuations manifest in observable quantities.
@@ Acknowledgements
We thank X. Turkeshi, B. Pain, and J. Kudler-Flam for their insightful discussions, M. Buchhold, L. Foini and Y. Fyodorov for valuable feedback on our work, and J. Kurchan for collaboration on related projects. We acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 -390534769, and DFG Collaborative Research Center (CRC) 183 Project No. 277101999 - project B02.
Appendix A Diagrammatic representation
of the Full Eigenstate Thermalization Hypothesis